# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v2

@@ 1.4 Significance
We then show how to modify this algorithm to compute the (monotone) Fréchet distance (in cases when the morphing was not already monotone), and how to make it handle large inputs quickly via simplifications (handling all the technical difficulties this gives rise to). Finally, we implemented our new algorithms in Julia and Python, and made them publicly available as standard packages. Making the computation of the Fréchet distance accessible to a wider audience.
@@ 3.1 Definition and basic algorithm
Figure 3.1: Left: Two curves. Right: Their elevation function, and the associated graph.
@@ 4.1 Direct Monotonization
Figure 4.1: Left: Two curves. Right: Their Fréchet morphing. Since the Fréchet morphing computed inside each cell of the free space diagram is a segment, it is enough to mark all the critical configurations when the morphing enters/leaves a cell. Each critical configuration is depicted on the right by a segment connecting the two points being matched. In between two such configurations, the morphing is the linear interpolation between the two configurations. See his link for an animation of this specific morphing.
@@
For many natural inputs $\omega\left(\mathcalb{m}_{\textsf{VE}}\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$. Specifically, this occurs if either $\mathcalb{m}_{\textsf{VE}}$ is monotone to begin with or if the “zigzagging” in $\mathcalb{m}_{\textsf{VE}}$ is happening in parts where the leash is relatively short, and thus the fixup due to direct monotonization does not affect the global bottleneck. Note that in this case, $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$, and so by Observation 4.3, $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)$ is an optimal morphing realizing $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$.
@@ 4.1.2 Monotonization via refinement
4.1.2.1 The lower/upper bound.
Let $\mathcalb{m}_{\textsf{VE}}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. By Observation 4.3, $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\in[\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$. Thus we have initial upper and lower bounds for $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and our approach is to repeatedly improve these bounds until they are equal, at which point the computed morphing realizes the Fréchet distance between the two curves.
Observation 4.4.
@@
Indeed, create $\pi^{\prime}$ and $\sigma^{\prime}$ by introducing vertices along $\pi$ and $\sigma$ respectively, such that no edge has length exceeding $\varepsilon/2$. Recall that by Observation 3.9, $\mathcalb{m}_{\textsf{VE}}^{\prime}$ is already monotone over vertices, and thus $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}^{\prime}\right)$ does not change the edges from the curves that points in the morphing are mapped to. Thus the direct monotonizations of the VE morphings of $\pi^{\prime}$ and $\sigma^{\prime}$ can increase the distance by at most twice the maximum edge length.
@@ 4.3.1 Combining morphings
4.3.1.1 Getting a monotone morphing via simplification.
@@
The algorithm for combining two morphings is a linear time algorithm similar somewhat in nature to merge (from merge-sort). We omit the straightforward but tedious details.
@@ 4.5.2 The sweep distance
4.5.2.1 The cost of morphing along a segment inside a cell.
@@
4.5.2.2 Computing the sweep distance.
@@ 4.6.1 Fast extractor
The curve $\pi$ is uniformly parameterized on the interval $I=[0,\left\|{\pi}\right\|]$, see Definition 1.1. The morphing $\mathcalb{m}^{\prime}$ matches $\pi(t)$ with the point $\Downarrow_{\tau}\!\!\left(\pi(t)\right)$. Let $f(t)=\left\|{\Downarrow_{\tau}\!\!\left(\pi(t)\right)-z}\right\|$ be the normalized parameterization of this projection point on $\tau$. Thus, for all $t$, $f(t)\in[0,\left\|{\tau}\right\|]$, $f(0)=0$, and $f(\left\|{\pi}\right\|)=\left\|{\tau}\right\|$. The max version of $f$ is $F(t)=\max_{x:[0,t]}f(x)$, and the morphing $\mathcalb{m}$, for all $t$, maps $\pi(t)$ to $\tau(F(t))$. A maximal interval $J=[\alpha,\beta]\subseteq[0,\left\|{\pi}\right\|]$, with $\alpha<\beta$, such that $F(J)$ is a single value, is a plateau. For a plateau the morphing $\mathcalb{m}$ matches all the points of $\pi(J)$ to the point $q=\tau(F(J))$. A point $x\in[0,\left\|{\pi}\right\|]$ not contained in any plateau, is matched by $\mathcalb{m}$, to the same point it was matched by $\mathcalb{m}^{\prime}$, that is $\tau(f(x))=\Downarrow_{\tau}\!\!\left(\pi(x)\right)$.
@@
4.6.1.3 Extracting the simplification (extract).
@@
5 Implementation and experiments
To demonstrate the practicality of our techniques, we have developed open-source packages that implement our algorithms. The technologies used were chosen for ease of implementation while still providing performance suitable for processing large inputs. In particular, the most expensive single operation performed by algorithms described in this work are the numerical distance computations between points and segments, so preference was given to technologies that could accelerate these computations.
We emphasize however that we have not done low-level optimization of our code for the sake of performance. For instance, better performance might follow by implementing our algorithm in C++ or using squared distances. Our optimizations tend to be more algorithmic in nature (as described in the rest of paper), by aggressively using simplification, and upper/lower bound computations to guarantee we computed the optimal Fréchet distance.
Importantly, because our algorithms use the retractable VE-Fréchet as our building block, the Fréchet morphings our implementation computes look better than what the standard Fréchet matching would give.
5.1 Datasets and hardware
The tests were all performed on a Linux system with 64GB of memory and an Intel i7-11700 CPU with 16 threads. This system is a fast desktop, but far from the best possible hardware currently available.
We use several real-world datasets featured in prior works to demonstrate the effectiveness of our algorithms. Specifically, we use curves taken from the GeoLife dataset zfxml-ggtdu-11 (as used in bkn-wdfpa-21), a stork migration dataset rktzw-eabgc-18, (retrieved from Movebank rktzw-deabg-18, used in bbknp-klmct-20), and a pigeon flight dataset pgrbvg-pigeon-17. In addition, we use datasets from Bringmann et al. bkn-wdfpa-21 when comparing our implementation to theirs.
5.2 Python
For the Python implementation, we represented the curves using NumPy arrays numpy-20, where each row in the array is a single point on the input curve. This representation is memory efficient and allows for fast computations of distances between points using library primitives. We also made heavy use of Numba lps-nlpjc-15 to further improve performance, as many of the algorithms discussed in this work are iterative and thus easy to accelerate using the library.
The results of running benchmarks on our Python implementation are shown in Figure 5.3. Of note is that the Python implementation is slower than the Julia implementation on the computation of the exact distance. This can likely be improved through the use of data structures that support more efficient refinement operations. We did not include a column for a 1.001 approximation, as our implementation encountered errors with numerical precision. The runtimes also do not include the Numba JIT compilation times.
5.3 Julia
The Julia code is about 5000 lines for the library, and additional 5000 for the examples code. It implements the algorithms described in this paper faithfully, with some additional hacks to handle floating point issues. Julia has (surprisingly easy to use) support for multi-threading (unlike C++ or Python), and we used it to speed up various operations.
The Julia package is available at https://github.com/sarielhp/FrechetDist.jl. Animations and examples computed by the Julia code are available at https://sarielhp.org/p/24/frechet_ve/examples/.
Input P # Q # $\approx$4 $\approx$1.1 $\approx$1.01 $\approx$1.001 Exact VER
1 10,406 11,821 0.106 0.256 4.164 16.639 0.476 27.743
2 16,324 14,725 0.114 0.428 0.985 4.762 3.522 19.734
3 22,316 4,613 0.151 0.651 1.207 3.262 0.882 7.138
4 56,730 91,743 0.578 1.207 4.215 28.109 4.259 ---
5 6,103 9,593 0.043 0.040 0.305 0.519 0.484 5.678
6 2,702 1,473 0.034 0.084 1.367 1.387 0.160 1.081
7 1,068 1,071 0.026 0.066 0.048 0.049 0.044 0.011
8 864 1,168 0.018 0.044 0.142 0.124 0.069 0.080
Figure 5.1: Julia performance on various inputs. All running times are is seconds. The ##\## columns specify the number of vertices in each input. The single missing running time is for a case where the program ran out of memory.
5.4 Experiments
5.4.1 Julia vs C++
We compared our implementation performance in Julia to the C++ implementation of Bringmann et al. bkn-wdfpa-21 which seems to be the state of the art. Bringmann et al. also provides code that downloads and generates Fréchet distance tests, which we used ourselves. Of note is that the algorithm of Bringmann et al. uses completely different techniques from our algorithms.
We used three data sets. Each dataset is composed of numerous input curves, and then a list of pairs of the curves together with a threshold, and the task at hand is to decide whether the Fréchet distance is above/below/equal to the threshold. See Figure 5.2. This task at hand is naturally friendly to simplifications – we implemented a fast simplification data-structure. We iteratively simplify the two curves being compared to progressively finer resolutions, comparing them using the (monotonized) VE-Fréchet distance. If the comparison is outside the error interval, the result is returned, otherwise, the algorithm goes to the next finer resolution.
The GeoLife dataset benefits tremendously from the simplification strategy, and (maybe surprisingly) our Julia program is four time faster than the C++ code, see Figure 5.2. The worst dataset for the Julia implementation is the characters dataset where it seems to be about twice slower than the C++ code. This dataset both does not benefit from simplification, and also seem to give rise to many non-monotone morphings, which makes our implementation slower.
The third dataset (Sigspatial) seems to also be somewhat resistant to approximation, and our code is slightly slower than the C++ code. One can do massive preprocessing and precompute a simplifications hierarchy for each input curves. This improves the query process, but makes the preprocessing more expensive. Thus, storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries.
As for the general C++ vs Julia performance comparison. This seems to be case dependent, with reported cases where either is faster. In our cases, since we did not implement our library in C++, we do not know the situation for certain. The significant speed improvement in the case of the GeoLife input suggests our algorithm is much faster in some cases where simplification helps. Note, that while our code in Julia uses a single thread, it uses automatic garbage collector that runs in a parallel thread.
Polygonal hierarchies and fast simplification.
A strategy that works quite well for this task is to precompute for all input curves the fast simplification extractor of Lemma 4.25. Then given a query one can simplify the two curves quickly to the right resolution $\delta$ (initially guessed to be some large value), using this extractor, then apply the algorithm of Lemma 4.27 to get a simplification with few vertices, now compute the Fréchet distance, and repeat the process, with smaller $\delta$, till the query is resolved. By caching the simplifications of the input curves, one can achieve further speedup.
Input # pairs C++ Julia Julia MT sigspatial $322,000$ $145$ $149$ $36$ Characters $253,000$ $90$ $183$ $45$ Geolife $322,000$ $646$ $136$ $63$
Figure 5.2: Julia/C++ comparison. The C++ implementation is from Bringmann et al. bkn-wdfpa-21. Running times are in seconds. Julia MT stands for the multi-threaded implementation – since multi-threading is easy in Julia, and the task at hand is easily parallizable, we tested this, but of course this should not be taken as a direct (or remotely fair) comparison to the C++ implementation, and is provided for the reader amusement. In particular, the multi-threading done is pretty naive, and better performance should be possible by better fine-grained partition of the tasks. The tests we performed on a Linux system with 64GB memory with Intel i7-11700 CPU, with a decent 16 threads CPU, but far from the fastest hardware currently available.
Input P # Q # $\approx$4 $\approx$1.1 $\approx$1.01 Exact VER
1 10,400 11,815 0.071 0.250 5.372 0.593 92.869
2 16,318 14,719 0.091 0.240 1.142 69.525 66.844
3 22,310 4,607 0.087 0.232 1.041 49.426 25.032
4 56,730 91,743 0.307 0.563 1.892 2.390 ---
5 6,103 9,593 0.042 0.032 0.175 3.657 19.618
6 2,696 1,467 0.077 0.294 4.913 2.946 283.538
7 1,062 1,065 0.145 0.275 0.455 0.666 382.455
8 858 1,162 0.034 0.100 1.537 0.159 0.709
Input Description
1 birds: 1787_1 / 1797_1
2 birds: 2307_3 / 2859_3
3 birds: 2322_2 / 1793_4
4 GeoLife 20080928160000 / 20081219114010
5 GeoLife 20090708221430 / 20090712044248
6 Pigeons RH887_1 / RH887_11
7 Pigeons C369_5 / C873_6
8 Pigeons C360_10 / C480_9
5.4.2 Julia vs Python
Figure 5.4 lists the datasets we used in comparing the two implementations. See Figure 5.3 and Figure 5.1 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it seemed to be significantly slower and numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster), but it is clear that the Python implementation is fast and robust enough to be used in practice.

### v3

@@ 1.4 Significance
We then show how to modify this algorithm to compute the (monotone) Fréchet distance (in cases when the morphing was not already monotone), and how to make it handle large inputs quickly via simplifications (handling all the technical difficulties this gives rise to).
Finally, we implemented our new algorithms in Julia and Python, and made them publicly available as standard packages, thus making the computation of the Fréchet distance accessible to a wider audience. Using such packages in Python/Julia is significantly easier than using any not pre-compiled code in C++.
Summary.
The combination of simplification, new (and not so new) algorithmic ideas (such as retractablity via Dijkstra, and VE-Fréchet, among others), and careful implementation, leads to fast performance in practice beating other implementations. We discuss our implementation next — the algorithmic ideas described above are covered in detail in the appendices.
2 Implementation and experiments
To demonstrate the practicality of our techniques, we have developed open-source packages that implement our algorithms. Our implementations follow the ideas described in the paper – aggressively using simplification, and upper/lower bound computations to guarantee we computed the optimal Fréchet distance. Furthermore, because our algorithms use the retractable VE-Fréchet as our building block, the Fréchet morphings our implementation computes look better than what the standard Fréchet algorithms would return.
To see numerous animations and more information, the reader is encouraged to visit the (anonymized) site: https://frechet.xyz.
Total RT in seconds
Dataset # pairs avg # points C++ Julia Julia MT C++/Julia
sigspatial $322,000$ $247.8$ $145$ $102$ $25$ $\mathbf{1.42}$
Characters $253,000$ $120.9$ $90$ $72$ $18$ $\mathbf{1.25}$
Geolife $322,000$ $1080.4$ $646$ $70$ $30$ $\mathbf{9.22}$
All tests combined 881 244 73 $\mathbf{3.61}$
Figure 2.1: Julia/C++ comparison, demonstrating that the Julia implementation is significantly faster. The C++ implementation is from Bringmann et al. bkn-wdfpa-21. Running times are in seconds. Julia MT stands for the multi-threaded implementation – since multi-threading is easy in Julia, and the task at hand is easily parallizable, we tested this, but of course this should not be taken as a direct (or remotely fair) comparison to the C++ implementation, and is provided for the reader amusement. In particular, the multi-threading done is pretty naive, and better performance should be possible by better fine-grained partition of the tasks. The tests were performed on a Linux system with 64GB memory with Intel i7-11700 CPU, with a decent 16 threads CPU, but far from the fastest hardware currently available. The “# pairs” column is the number of pairs of curves that had their Fréchet distance compared during this test. The “avg # points” is the average number of vertices per curve in this input set – underlying the benefit of simplification for the GeoLife input set.
2.1 Hardware and Implementations
The tests were all performed on a Linux system with 64GB of memory and an Intel i7-11700 CPU with 16 threads. This system is a fast desktop, but far from the best possible hardware currently available. We were also able to perform development for the experiments on a laptop with substantially weaker hardware, giving anecdotal evidence of the scalability of our algorithms. We use several real-world datasets featured in prior works to demonstrate the effectiveness of our algorithm.
Here, we discuss the individual implementations used in our experiments. The first two (Julia and Python) use the algorithmic techniques discussed in this work. The third is by Bringmann et al. bkn-wdfpa-21 written in C++, which uses completely different algorithmic ideas. We refer to each implementation by language for convenience, although our tasks are meant to evaluate effectiveness of the algorithmic approaches.
2.1.1 Julia
The Julia code is about 5000 lines for the library, and an additional 5000 for the examples/testing code. It implements the algorithms described here faithfully, with some additional hacks to handle floating point issues. Julia has (surprisingly easy to use) support for multi-threading (unlike C++ or Python), and we tested also a multi-threaded version (this is shown in the MT column in Figure 2.1). The anonymized code is available from https://frechet.xyz/, including instructions how to replicate our tests.
The Julia package is available at https://github.com/sarielhp/FrechetDist.jl. Animations and examples computed by the Julia code are available at https://frechet.xyz/.
Julia low-level optimizations.
We were able to significantly improve the performance of the Julia code using some standard low-level optimizations: (i) Julia performs inlining code automatically, but explicitly forcing it to inline some low-level key functions (such as computing the distance between two points), improved the performance measurably ($>10\%$). Julia also has high overhead for creating structs that get destroyed immediately (essentially because it allocates them on the heap, instead of the stack like C++). Thus, by rewriting the code so that computing the distance of a point from a segment is done directly on the endpoints of the segment (where the segment is an edge of a polygon) also improved performance. Caching simplifications, so that one can extract/compute new simplification quickly, among other ideas, see Section 5.6, were useful in creating fast code.
One allocation to rule them all.
Since Julia performance suffers when a lot of dynamic allocations are done (because of the automatic garbage collection), it turned out that implementing the retractable VE-Fréchet without using hashing (for the vertices of the graph) resulted in much faster code. To this end, one pre-allocates a quadratic size array representing the vertices of the grid (but exploring only the relevant portions of the grid/this array) — this works because the allocation/initialization of this (single!) array is extremely fast. This seems to yields significant speedups somewhere up to ten times faster than the hashing version, depending on the input (for large inputs, the two implementations seem to have similar running times).
2.1.2 Python
For the Python implementation, we represented curves using NumPy arrays numpy-20, where each row in the array is a single point on the input curve. This representation is memory efficient and allows for fast computations of distances between points using library primitives. We also made heavy use of Numba lps-nlpjc-15 to further improve performance, as many of the algorithms discussed in this work are iterative and thus easy to accelerate using the library.
Numerical Issues.
Although not exclusive to this implementation, the Python code suffers from some issues related to numerical precision. Specifically, this implementation struggles with underflow of floating point numbers, which seems to occur in situations where consecutive points are too close in the input data (i.e., the distance between points is too small). However, this doesn’t represent a meaningful distance between these points, and this can be resolved by adding simplification as a preprocessing step (such as the algorithm in Section 5.6).
Our experiments are meant to emphasize the performance of our algorithmic techniques, and we do not have any tasks specifically designed to evaluate issues with floating point numbers or methods which can be used to combat this problem. In particular, we did not encounter these issues in any significance with the datasets mentioned in this work. This is a potential avenue for future work if given data that encounters these issues frequently.
2.1.3 C++
The C++ implementation is provided by Bringmann et al. bkn-wdfpa-21, which seems to be the state of the art (they claim to be faster than any previously available implementation). The algorithm uses completely different algorithmic techniques to those of this work, providing a point of comparison for the scalability of our implementations.
Informally, the implementation of Bringmann et al. bkn-wdfpa-21 builds a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
Note, that unlike Bringmann et al. bkn-wdfpa-21, our implementation can compute the exact Fréchet distance — although in some cases, because of numerical issues the “exact” results are just (high quality) approximations, see Remark 1.6.
2.2 Experiments
To compare the implementations, we performed benchmarks using two different tasks using real-world datasets. These were selected to determine the scalability of the algorithms relative to the large size of real-world data. This section is organized by the different tasks performed.
2.2.1 Direct Fréchet distance computations
To compare the Julia and Python implementations of our algorithms, we directly computed the Fréchet distance at different levels of approximation using our algorithms.
This was done on individual pairs of curves selected from different datasets. Specifically, we use curves taken from the GeoLife dataset zfxml-ggtdu-11 (as used in bkn-wdfpa-21), a stork migration dataset rktzw-eabgc-18, (retrieved from Movebank rktzw-deabg-18, used in bbknp-klmct-20), and a pigeon flight dataset pgrbvg-pigeon-17. Figure 2.2 lists the individual curves from each dataset used in our benchmarks.
See Figure 2.4 and Figure 2.3 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it was numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster), but it is clear that the Python implementation is fast and robust enough to be used in practice. The runtimes also do not include the Numba JIT compilation times.
Specifically, the slow performance observed on some test cases for the Python implementation may be caused by using sub-optimal data structures for some simplification tasks. We hope to improve this performance in further development of this implementation.
2.2.2 Fréchet Decider
We compared our implementation performance in Julia to the C++ implementation. We evaluated our implementations with the same benchmarking task as Bringmann et al. bkn-wdfpa-21, using the provided code that downloads and generates Fréchet distance tests. We emphasize that this task and datasets were introduced prior to to this work, and were not selected to be favorable to the techniques introduced here.
This task used curves from three datasets; the dataset from the SIGSPATIAL GIS Cup wo-asgcr-2018, a dataset of handwritten characters (Characters) w-ct-08, and the GeoLife dataset zfxml-ggtdu-11. Each dataset is composed of numerous input curves, and then a list of pairs of curves together with a threshold, and the goal is to decide whether the Fréchet distance is above/below/equal to this threshold. See Figure 2.1. This task is naturally friendly to simplifications – we implemented a fast simplification data-structure. We iteratively simplify the two curves being compared to progressively finer resolutions, comparing them using the (monotonized) VE-Fréchet distance. If the comparison is outside the error interval, the result is returned, otherwise, the algorithm goes to the next finer resolution.
The GeoLife zfxml-ggtdu-11 dataset benefits tremendously from the simplification strategy, and the Julia implementation is more than $9$ times faster than the C++ code, see Figure 2.1. The worst dataset for the Julia implementation is the characters dataset w-ct-08 where it seems to be about $25\%$ faster than the C++ code. This dataset both does not benefit from simplification, and also seems to give rise to many non-monotone morphings. The Sigspatial seems be somewhat resistant to approximation, and our code is only about $40\%$ faster than the C++ code in this case. We emphasize that each test in Figure 2.1 involves the execution of well over $>250,000$ runs of the algorithm for different curves, so these results seems quite representative of real world performance.
As for the general C++ vs Julia performance comparison, this seems to be case dependent, with reported cases where either is faster. The common belief is that a careful enough implementation in C++ would be faster from Julia implementation (i.e., some of Julia libraries are implemented in C++).
In particular, the significant speed (646 seconds vs 70) improvement in the case of the GeoLife input suggests our algorithm is much faster in cases of longer curves where simplification really helps. Note that while our code in Julia uses a single thread, it uses an automatic garbage collector that runs in parallel threads.
2.3 Discussion
Overall, it is clear that our Julia implementation is faster than other currently available implementations for computing the Fréchet distance. Beyond that Julia seems like a great programming language to develop geometric algorithms — providing a combination of a safe, multi-threaded, high-level programming language with the performance of C++, without the pain involved in working in C++.
One optimization used by the Julia code, that should be generally useful, is the following. As a preprocessing step, precompute a hierarchy of simplified curves for each input curve. This improves the query process, but makes the preprocessing more expensive. Thus, storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries. We emphasize that the reported running times include this (light) preprocessing stage (interestingly, the SIGSPTIAL wo-asgcr-2018 competition allowed such preprocessing to not be included in the overall running time).
Input Description
1 birds: 1787_1 / 1797_1
2 birds: 2307_3 / 2859_3
3 birds: 2322_2 / 1793_4
4 GeoLife 20080928160000 / 20081219114010
5 GeoLife 20090708221430 / 20090712044248
6 Pigeons RH887_1 / RH887_11
7 Pigeons C369_5 / C873_6
8 Pigeons C360_10 / C480_9
Input P # Q # $\approx$4 $\approx$1.1 $\approx$1.01 $\approx$1.001 Exact VER
1 10,406 11,821 0.106 0.256 4.164 16.639 0.476 27.743
2 16,324 14,725 0.114 0.428 0.985 4.762 3.522 19.734
3 22,316 4,613 0.151 0.651 1.207 3.262 0.882 7.138
4 56,730 91,743 0.578 1.207 4.215 28.109 4.259 ---
5 6,103 9,593 0.043 0.040 0.305 0.519 0.484 5.678
6 2,702 1,473 0.034 0.084 1.367 1.387 0.160 1.081
7 1,068 1,071 0.026 0.066 0.048 0.049 0.044 0.011
8 864 1,168 0.018 0.044 0.142 0.124 0.069 0.080
Figure 2.3: Julia performance on various inputs. All running times are is seconds. The ##\## columns specify the number of vertices in each input. The single missing running time is for a case where the program ran out of memory. The quality of approximation is in the title of the column. The VER column is for running the VE-Fréchet (retractable) algorithm on the original input curves, which is much slower than the exact algorithm, which computes the exact Fréchet distance, but uses simplification internally for speed.
Input P # Q # $\approx$4 $\approx$1.1 $\approx$1.01 Exact VER
1 10,400 11,815 0.071 0.250 5.372 0.593 92.869
2 16,318 14,719 0.091 0.240 1.142 69.525 66.844
3 22,310 4,607 0.087 0.232 1.041 49.426 25.032
4 56,730 91,743 0.307 0.563 1.892 2.390 ---
5 6,103 9,593 0.042 0.032 0.175 3.657 19.618
6 2,696 1,467 0.077 0.294 4.913 2.946 283.538
7 1,062 1,065 0.145 0.275 0.455 0.666 382.455
8 858 1,162 0.034 0.100 1.537 0.159 0.709
@@ 4.1 Definition and basic algorithm
Figure 4.1: Left: Two curves. Right: Their elevation function, and the associated graph. See here for more info.
@@ 5.1 Direct Monotonization
Figure 5.1: Left: Two curves. Right: Their Fréchet morphing. Since the Fréchet morphing computed inside each cell of the free space diagram is a segment, it is enough to mark all the critical configurations when the morphing enters/leaves a cell. Each critical configuration is depicted on the right by a segment connecting the two points being matched. In between two such configurations, the morphing is the linear interpolation between the two configurations. See here for an animation of this specific morphing.
@@
For many natural inputs $\omega\left(\mathcalb{m}_{\textsf{VE}}\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$. Specifically, this occurs if either $\mathcalb{m}_{\textsf{VE}}$ is monotone to begin with or if the “zigzagging” in $\mathcalb{m}_{\textsf{VE}}$ is happening in parts where the leash is relatively short, and thus the fixup due to direct monotonization does not affect the global bottleneck. Note that in this case, $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$, and so by Observation 5.3, $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)$ is an optimal morphing realizing $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$.
@@ 5.1.2 Monotonization via refinement
The lower/upper bound.
Let $\mathcalb{m}_{\textsf{VE}}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. By Observation 5.3, $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\in[\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$. Thus we have initial upper and lower bounds for $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and our approach is to repeatedly improve these bounds until they are equal, at which point the computed morphing realizes the Fréchet distance between the two curves.
Observation 5.4.
@@
Indeed, create $\pi^{\prime}$ and $\sigma^{\prime}$ by introducing vertices along $\pi$ and $\sigma$ respectively, such that no edge has length exceeding $\varepsilon/2$. Recall that by Observation 4.9, $\mathcalb{m}_{\textsf{VE}}^{\prime}$ is already monotone over vertices, and thus $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}^{\prime}\right)$ does not change the edges from the curves that points in the morphing are mapped to. Thus the direct monotonizations of the VE morphings of $\pi^{\prime}$ and $\sigma^{\prime}$ can increase the distance by at most twice the maximum edge length.
@@ 5.3.1 Combining morphings
Getting a monotone morphing via simplification.
@@
The algorithm for combining two morphings is a linear time algorithm somewhat similar in nature to the merge subroutine used in merge-sort. We omit the straightforward but tedious details.
@@ 5.5.2 The sweep distance
The cost of morphing along a segment inside a cell.
@@
Computing the sweep distance.
@@ 5.6.1 Fast extractor
The curve $\pi$ is uniformly parameterized on the interval $I=[0,\left\|{\pi}\right\|]$, see Definition 1.1. The morphing $\mathcalb{m}^{\prime}$ matches $\pi(t)$ with the point $\Downarrow_{\tau}\!\!\left(\pi(t)\right)$.
Let $f(t)=\left\|{\Downarrow_{\tau}\!\!\left(\pi(t)\right)-z}\right\|$ be the normalized parameterization of this projection point on $\tau$. Thus, for all $t$, $f(t)\in[0,\left\|{\tau}\right\|]$, $f(0)=0$, and $f(\left\|{\pi}\right\|)=\left\|{\tau}\right\|$. The max version of $f$ is $F(t)=\max_{x:[0,t]}f(x)$, and the morphing $\mathcalb{m}$, for all $t$, maps $\pi(t)$ to $\tau(F(t))$. A maximal interval $J=[\alpha,\beta]\subseteq[0,\left\|{\pi}\right\|]$, with $\alpha<\beta$, such that $F(J)$ is a single value, is a plateau. For a plateau the morphing $\mathcalb{m}$ matches all the points of $\pi(J)$ to the point $q=\tau(F(J))$. A point $x\in[0,\left\|{\pi}\right\|]$ not contained in any plateau, is matched by $\mathcalb{m}$, to the same point it was matched by $\mathcalb{m}^{\prime}$, that is $\tau(f(x))=\Downarrow_{\tau}\!\!\left(\pi(x)\right)$.
@@
Extracting the simplification (extract).